# Localization of Dirac modes in the $\mathrm{SU}(2)$-Higgs model at finite temperature

### 19 Feb 2024 v2

@@ Abstract
We investigate the connection between localization of low-lying Dirac modes and Polyakov-loop ordering in the lattice $\mathrm{SU}(2)$-Higgs model at finite temperature, probed with static external staggered fermions. After mapping out the phase diagram of the model at a fixed temporal extension in lattice units, we study the localization properties of the low-lying modes of the staggered Dirac operator, how these properties change across the various transitions, and how these modes correlate with the gauge and Higgs fields. We find localized low modes in the deconfined and in the Higgs phase, where the Polyakov loop is strongly ordered, but in both cases they disappear as one crosses over to the confined phase. Our findings confirm the general expectations of the “sea/islands” picture, and the more detailed expectations of its refined version concerning the favorable locations of localized modes, also in the presence of dynamical scalar matter.
@@ I Introduction
A model allowing one to carry out both these tests at once is the lattice fixed-length $\mathrm{SU}(2)$-Higgs model [36]. At zero temperature the phase diagram of this model has been studied in depth both with analytical [36] and numerical [37, 38, 39, 40, 41, 42] methods. This model has two parameters, namely the (inverse) gauge coupling $\beta$ and the Higgs-gauge coupling $\kappa$, and it displays two lines of transitions in the $(\beta,\kappa)$ plane as follows [42]:
@@
While fermions are absent in the $\mathrm{SU}(2)$-Higgs model, one can still probe this system using static external fermions coupled to the $\mathrm{SU}(2)$ gauge field, as pointed out above. One can then study how the corresponding Dirac spectrum behaves, and check what happens to the localization properties of its low modes across the various transitions, in particular as one crosses over to the Higgs phase starting from either the confined or the deconfined phase. Since eigenvalues and eigenvectors of the Dirac operator are nonlocal functions of the gauge fields, they can display non-analytic behavior even in the strip of the $(\beta,\kappa)$ plane where all local correlators are analytic functions of the couplings, and so they could allow one to sharply distinguish the confined and the Higgs phase. (A different approach to this issue, based on the analogies between gauge-Higgs theories and spin glasses, is discussed in the review Ref. [44] and references therein.)
In this paper we study the spectrum and the eigenvectors of the staggered lattice Dirac operator in the $\mathrm{SU}(2)$-Higgs model at finite temperature. After briefly describing the model, in section II we introduce the tools we use to investigate the localization properties of staggered eigenmodes. In section III we map out the phase diagram of the model at finite temperature, working at fixed temporal extension in lattice units. In section IV we analyze the staggered eigenmodes, focussing in particular on how their localization properties change across the transitions between the confined, deconfined, and Higgs phases. We then study in detail the correlation between eigenmodes and the gauge and Higgs fields, to identify the field fluctuations mostly responsible for localization. Finally, in section V we draw our conclusions and show some prospects for the future.
II $\mathrm{SU}(2)$-Higgs model and localization
In this section we describe the fixed-length $\mathrm{SU}(2)$-Higgs model, and discuss how to characterize the localization properties of Dirac modes, and how these correlate with the gauge and Higgs fields.
II.1 $\mathrm{SU}(2)$-Higgs model on the lattice
We study the lattice $\mathrm{SU}(2)$-Higgs model in 3+1 dimensions, defined by the action
@@ II.2 Localization of staggered eigenmodes
One can give up some of the accuracy but save a lot in computing effort by using the critical value of the spectral statistic, expected to be universal, to determine the mobility edge simply by looking for the point where the curve for $I_{s_{0}}$ crosses its critical value, $I_{s_{0},c}$ (see, e.g, Refs. [28, 23, 25, 29]). This critical value is not known for the symplectic class, but it can be determined by identifying the scale-invariant point in the spectrum at some point in the parameter space of the model under study (if one can find an Anderson transition, of course); the corresponding critical value can then be used in the rest of the analysis. Notice that one could estimate the mobility edge in a finite volume as the point where $I_{s_{0}}$ takes any chosen value intermediate between the RMT and the Poisson predictions, and this would converge to the correct value in the infinite-volume limit. In this respect, the choice of $I_{s_{0},c}$ is only the most convenient, as it is expected to minimize the magnitude of finite-size effects.
@@ III Phase diagram at finite temperature
A sketch of the resulting phase diagram is shown in Fig. 3, obtained by merging the various transition lines, defined by the peaks of the suceptibilities. The dashed line at low $\beta$ and large $\kappa$ signals a sizeable reduction in the strength of the transition there, as shown by all three observables. Except in this region, where they slighly deviate from each other, the transition lines between confined and Higgs phase obtained from the three different susceptibilities agree with each other, so we drew a single line. Similarly, the transition lines between confined and deconfined phase obtained from the plaquette and the Polyakov loop susceptibility agree with each other, so we drew a single line in this case as well.
@@ IV.3 $\beta$ and $\kappa$ dependence of the mobility edge
Figure 18: Mode size averaged up to $\lambda_{c}/2$, Eq. 37, at $\kappa=0.3$ in the deconfined phase (top panel) and at $\kappa=1.0$ in the Higgs phase (center panel), as a function of $\beta$, and at $\beta=2.6$ across the transition from the deconfined to the Higgs phase (bottom panel), as a function of $\kappa$. In all plots $N_{s}=20$ and $N_{t}=4$. The solid line in the top and center panels is a fit with a power-law function. The vertical and dashed lines in the bottom panel mark the critical value $\kappa_{\mathrm{loc}}$ and the corresponding error band [see after Eq. (36)].
@@
yield $\beta_{\mathrm{loc}}=2.3318(15)$ in the deconfined phase, and $\beta_{\mathrm{loc}}=2.0499(92)$ in the Higgs phase, both in the crossover region, and in reasonable agreement with the determinations based on the extrapolation of the mobility edge. Here one should take into account that the functional form Eq. (38) is not fully justified, as the mode size cannot diverge in a finite volume, and there is no reason to assume that the mode size goes to zero at large $\beta$. Nonetheless, one obtains decent fits (see the resulting fit parameters and $\chi^{2}$ in Tab. 5); adding a costant term makes them worse. On top of this, the error estimates do not include any uncertainty due to finite-size effects, which are large near $\beta_{\mathrm{loc}}$. For completeness, in the bottom panel of Fig. 18 we show $N_{t}V\cdot\langle\mathrm{PR}\rangle_{\lambda<\frac{\lambda_{c}}{2}}$ as a function of $\kappa$ at constant $\beta=2.6$ across the two phases. Here the data indicate a finite mode size at all $\kappa$, with a change from a constant to a steadily decreasing trend taking place at the crossover between the deconfined and the Higgs phase, showing that localized modes shrink rapidly as one moves deeper in the Higgs phase and the Polyakov-loop expectation value increases (see Fig. 1).
@@ IV.4 Correlation with bosonic observables and sea/island mechanism
We now proceed to discuss our results on the correlation of staggered eigenmodes with the gauge and Higgs fields. To this end, the most informative quantities are the centered and normalized observables $\widehat{\mathcal{U}}$, $\widehat{\mathcal{P}}$ and $\widehat{\mathcal{G}}$, defined in Eq. (21), that take into account the width of the distribution of the relevant bosonic observables. Our results for these quantities are shown in Figs. 1921. The statistical error on the numerical estimate of these quantities is obtained by first determining the jackknife error on $\mathcal{U}$, $\mathcal{P}$ and $\mathcal{G}$, and correspondingly on $\langle U\rangle$, $\langle P\rangle$, $\langle G\rangle$ and on $(\delta U)^{2}$, $(\delta P)^{2}$, $(\delta G)^{2}$, followed by linear error propagation. Correlations with Polyakov-loop and plaquette fluctuations are always negative, showing that low modes prefer locations where these quantities fluctuate to values below their average. Correlations with gauge-Higgs coupling term fluctuations are again negative in the confined and in the Higgs phase, while they are essentially compatible with zero in the deconfined phase.
@@ V Conclusions
A strong connection has emerged in recent years between the deconfinement phase transition in gauge theories with or without fermionic matter, and the change in the localization properties of low Dirac modes [3, 4, 5, 6, 7, 8, 9, 10, 18, 19, 20, 21, 22, 12, 13, 14, 15, 23, 28, 24, 25, 26, 29, 27, 16]. In this paper we extended this line of research by studying the lattice $\mathrm{SU}(2)$-Higgs model with a Higgs field of fixed length [36, 37, 38, 39, 40, 41, 42] at finite temperature, probed with external static fermions. The extension is twofold. On the one hand, this model has dynamical scalar rather than fermionic matter: while one still expects localized modes in the deconfined phase of the model, as the nature of the dynamical matter does not affect the general argument for localization [12, 13, 14, 15, 10, 16], it is nonetheless useful to verify this explicitly. On the other hand, and more interestingly, the two-parameter phase diagram of this model displays a third phase besides the confined and deconfined phases, i.e., the Higgs phase: one can then check whether or not modes are localized in this phase, and if so whether the onset of localization is related in any way to the thermodynamic transition.
@@
By means of numerical simulations, we have demonstrated that localized modes are indeed present both in the deconfined and in the Higgs phase. In both cases, the mobility edge separating localized and delocalized modes in the spectrum decreases as one moves towards the confined phase, and disappears as one reaches the crossover region. At the transition between the deconfined and the Higgs phase, instead, the dependence of the mobility edge on the gauge-Higgs coupling constant changes from almost constant to steadily increasing. These findings provide further support to the universal nature of the sea/islands picture of localization [12, 13, 14, 15, 10, 16] in a previously unexplored setup in the presence of dynamical scalar matter.
@@
A possible extension of this work would be a study of the low $\beta$, large $\kappa$ corner of the phase diagram, where the crossover becomes very weak, in order to check if the line of “geometric” transitions where the mobility edge in the Dirac spectrum vanishes extends all the way to $\beta=0$, or if instead it has an endpoint. This is interesting also in connection with the “spin glass” approach of Ref. [44]: since in that region of parameter space this predicts a transition line clearly distinct from the one found with more traditional approaches based on gauge fixing, one would like to compare this line with the one defined by the vanishing of the mobility edge (if the latter exists). A different direction would be the study of the localization properties of the eigenmodes of the covariant Laplacian, extending to finite temperature and dynamical scalar matter the work of Refs. [62, 63].
@@ Acknowledgments
We thank T.G. Kovács for useful discussions and a careful reading of the manuscript. MG was partially supported by the NKFIH grant KKP-126769.

### 12 Mar 2024 v3

# Localization of Dirac modes in the $\mathrm{SU}(2)$ Higgs model at finite temperature

@@ Abstract
We investigate the connection between localization of low-lying Dirac modes and Polyakov-loop ordering in the lattice $\mathrm{SU}(2)$ Higgs model at finite temperature, probed with the staggered Dirac operator. After mapping out the phase diagram of the model at a fixed temporal extension in lattice units, we study the localization properties of the low-lying modes of the staggered Dirac operator, how these properties change across the various transitions, and how these modes correlate with the gauge and Higgs fields. We find localized low modes in the deconfined and in the Higgs phase, where the Polyakov loop is strongly ordered, but in both cases they disappear as one crosses over to the confined phase. Our findings confirm the general expectations of the “sea/islands” picture, and the more detailed expectations of its refined version concerning the favorable locations of localized modes, also in the presence of dynamical scalar matter.
@@ I Introduction
A model allowing one to carry out both these tests at once is the lattice fixed-length $\mathrm{SU}(2)$ Higgs model [36]. At zero temperature the phase diagram of this model has been studied in depth both with analytical [36] and numerical [37, 38, 39, 40, 41, 42] methods. This model has two parameters, namely the (inverse) gauge coupling $\beta$ and the Higgs-gauge coupling $\kappa$, and it displays two lines of transitions in the $(\beta,\kappa)$ plane as follows [42]:
@@
While fermions are absent in the $\mathrm{SU}(2)$ Higgs model, one can still probe this system using static external fermions coupled to the $\mathrm{SU}(2)$ gauge field, as pointed out above. One can then study how the corresponding Dirac spectrum behaves, and check what happens to the localization properties of its low modes across the various transitions, in particular as one crosses over to the Higgs phase starting from either the confined or the deconfined phase. Since eigenvalues and eigenvectors of the Dirac operator are nonlocal functions of the gauge fields, they can display non-analytic behavior even in the strip of the $(\beta,\kappa)$ plane where all local correlators are analytic functions of the couplings, and so they could allow one to sharply distinguish the confined and the Higgs phase. (A different approach to this issue, based on the analogies between gauge-Higgs theories and spin glasses, is discussed in the review Ref. [44] and references therein.)
In this paper we study the spectrum and the eigenvectors of the staggered lattice Dirac operator in the $\mathrm{SU}(2)$ Higgs model at finite temperature. After briefly describing the model, in section II we introduce the tools we use to investigate the localization properties of staggered eigenmodes. In section III we map out the phase diagram of the model at finite temperature, working at fixed temporal extension in lattice units. In section IV we analyze the staggered eigenmodes, focussing in particular on how their localization properties change across the transitions between the confined, deconfined, and Higgs phases. We then study in detail the correlation between eigenmodes and the gauge and Higgs fields, to identify the field fluctuations mostly responsible for localization. Finally, in section V we draw our conclusions and show some prospects for the future.
II $\mathrm{SU}(2)$ Higgs model and localization
In this section we describe the fixed-length $\mathrm{SU}(2)$ Higgs model, and discuss how to characterize the localization properties of Dirac modes, and how these correlate with the gauge and Higgs fields.
II.1 $\mathrm{SU}(2)$ Higgs model on the lattice
We study the lattice $\mathrm{SU}(2)$ Higgs model in 3+1 dimensions, defined by the action
@@ II.2 Localization of staggered eigenmodes
One can give up some of the accuracy but save a lot in computing effort by using the critical value of the spectral statistic, expected to be universal, to determine the mobility edge simply by looking for the point where the curve for $I_{s_{0}}$ crosses its critical value, $I_{s_{0},c}$ (see, e.g., Refs. [28, 23, 25, 29]). This critical value is not known for the symplectic class, but it can be determined by identifying the scale-invariant point in the spectrum at some point in the parameter space of the model under study (if one can find an Anderson transition, of course); the corresponding critical value can then be used in the rest of the analysis. Notice that one could estimate the mobility edge in a finite volume as the point where $I_{s_{0}}$ takes any chosen value intermediate between the RMT and the Poisson predictions, and this would converge to the correct value in the infinite-volume limit. In this respect, the choice of $I_{s_{0},c}$ is only the most convenient, as it is expected to minimize the magnitude of finite-size effects.
@@ III Phase diagram at finite temperature
A sketch of the resulting phase diagram is shown in Fig. 3, obtained by merging the various transition lines, defined by the peaks of the suceptibilities. The dashed line at low $\beta$ and large $\kappa$ signals a sizeable reduction in the strength of the transition there, as shown by all three observables. Except in this region, where they slightly deviate from each other, the transition lines between confined and Higgs phase obtained from the three different susceptibilities agree with each other, so we drew a single line. Similarly, the transition lines between confined and deconfined phase obtained from the plaquette and the Polyakov loop susceptibility agree with each other, so we drew a single line in this case as well.
@@ IV.3 $\beta$ and $\kappa$ dependence of the mobility edge
Instead of extrapolating in $\beta$, one could in principle explore the crossover region directly without particular problems, as there is no critical slowing down taking place there. However, our extrapolations convincingly show that $\lambda_{c}$ will be close to zero near the crossover to the confined phase. In the near-zero region the spectral density is low, and the effects of the approximate taste symmetry of staggered fermions on the spectrum is prominent. These effects consist in the formation of nearly degenerate multiplets of eigenvalues that distort the spectral statistics away from the expected universal behavior (see Ref. [4]), making our method unreliable. To cure this problem one needs to make the lattice volume large enough, so that the size of the would-be multiplets (which is controlled by the lattice spacing) becomes larger than the typical level spacing (which is controlled by the inverse lattice volume), and the approximate taste symmetry does not affect the short-range spectral statistics. This is numerically expensive, and we have preferred to adopt here a computationally less intensive method, leaving the direct investigation of the crossover region to future work.
@@
Figure 18: Mode size averaged up to $\lambda_{c}/2$, Eq. (37), at $\kappa=0.3$ in the deconfined phase (top panel) and at $\kappa=1.0$ in the Higgs phase (center panel), as a function of $\beta$, and at $\beta=2.6$ across the transition from the deconfined to the Higgs phase (bottom panel), as a function of $\kappa$. In all plots $N_{s}=20$ and $N_{t}=4$. The solid line in the top and center panels is a fit with a power-law function. The vertical and dashed lines in the bottom panel mark the critical value $\kappa_{\mathrm{loc}}$ and the corresponding error band [see after Eq. (36)].
@@
yield $\beta_{\mathrm{loc}}=2.3318(15)$ in the deconfined phase, and $\beta_{\mathrm{loc}}=2.0499(92)$ in the Higgs phase, both in the crossover region, and in reasonable agreement with the determinations based on the extrapolation of the mobility edge. Here one should take into account that the functional form Eq. (38) is not fully justified, as the mode size cannot diverge in a finite volume, and there is no reason to assume that the mode size goes to zero at large $\beta$. Nonetheless, one obtains decent fits (see the resulting fit parameters and $\chi^{2}$ in Tab. 5); adding a constant term makes them worse. On top of this, the error estimates do not include any uncertainty due to finite-size effects, which are large near $\beta_{\mathrm{loc}}$. For completeness, in the bottom panel of Fig. 18 we show $N_{t}V\cdot\langle\mathrm{PR}\rangle_{\lambda<\frac{\lambda_{c}}{2}}$ as a function of $\kappa$ at constant $\beta=2.6$ across the two phases. Here the data indicate a finite mode size at all $\kappa$, with a change from a constant to a steadily decreasing trend taking place at the crossover between the deconfined and the Higgs phase, showing that localized modes shrink rapidly as one moves deeper in the Higgs phase and the Polyakov-loop expectation value increases (see Fig. 1).
@@ IV.4 Correlation with bosonic observables and sea/island mechanism
We now proceed to discuss our results on the correlation of staggered eigenmodes with the gauge and Higgs fields. To this end, the most informative quantities are the centered and normalized observables $\widehat{\mathcal{U}}$, $\widehat{\mathcal{P}}$, and $\widehat{\mathcal{G}}$, defined in Eq. (21), that take into account the width of the distribution of the relevant bosonic observables. Our results for these quantities are shown in Figs. 1921. The statistical error on the numerical estimate of these quantities is obtained by first determining the jackknife error on $\mathcal{U}$, $\mathcal{P}$, and $\mathcal{G}$, and correspondingly on $\langle U\rangle$, $\langle P\rangle$, $\langle G\rangle$ and on $(\delta U)^{2}$, $(\delta P)^{2}$, $(\delta G)^{2}$, followed by linear error propagation. Correlations with Polyakov-loop and plaquette fluctuations are always negative, showing that low modes prefer locations where these quantities fluctuate to values below their average. Correlations with gauge-Higgs coupling term fluctuations are again negative in the confined and in the Higgs phase, while they are essentially compatible with zero in the deconfined phase.
@@ V Conclusions
A strong connection has emerged in recent years between the deconfinement phase transition in gauge theories with or without fermionic matter, and the change in the localization properties of low Dirac modes [3, 4, 5, 6, 7, 8, 9, 10, 18, 19, 20, 21, 22, 12, 13, 14, 15, 23, 28, 24, 25, 26, 29, 27, 16]. In this paper we extended this line of research by studying the lattice $\mathrm{SU}(2)$ Higgs model with a Higgs field of fixed length [36, 37, 38, 39, 40, 41, 42] at finite temperature, probed with external static fermions. The extension is twofold. On the one hand, this model has dynamical scalar rather than fermionic matter: while one still expects localized modes in the deconfined phase of the model, as the nature of the dynamical matter does not affect the general argument for localization [12, 13, 14, 15, 10, 16], it is nonetheless useful to verify this explicitly. On the other hand, and more interestingly, the two-parameter phase diagram of this model displays a third phase besides the confined and deconfined phases, i.e., the Higgs phase: one can then check whether or not modes are localized in this phase, and if so whether the onset of localization is related in any way to the thermodynamic transition.
@@
By means of numerical simulations, we have demonstrated that localized modes are indeed present both in the deconfined and in the Higgs phase. In both cases, the mobility edge separating localized and delocalized modes in the spectrum decreases as one moves towards the confined phase, and extrapolates to zero as one reaches the crossover region. At the transition between the deconfined and the Higgs phase, instead, the dependence of the mobility edge on the gauge-Higgs coupling constant changes from almost constant to steadily increasing. These findings provide further support to the universal nature of the sea/islands picture of localization [12, 13, 14, 15, 10, 16] in a previously unexplored setup in the presence of dynamical scalar matter.
@@
A straightforward extension of this work would be the direct study of the region near the crossover to the confined phase, where larger volumes are required to apply our method than the ones employed here, in order to avoid the distortion effects of the approximate taste symmetry of staggered fermions on the spectral statistics [4]. Another possible extension would be a study of the low $\beta$, large $\kappa$ corner of the phase diagram, where the crossover becomes very weak, in order to check if the line of “geometric” transitions where the mobility edge in the Dirac spectrum vanishes extends all the way to $\beta=0$, or if instead it has an endpoint. This is interesting also in connection with the “spin glass” approach of Ref. [44]: since in that region of parameter space this predicts a transition line clearly distinct from the one found with more traditional approaches based on gauge fixing, one would like to compare this line with the one defined by the vanishing of the mobility edge (if the latter exists). A different direction would be the study of the localization properties of the eigenmodes of the covariant Laplacian, extending to finite temperature and dynamical scalar matter the work of Refs. [62, 63].
@@ Acknowledgments
We thank T. G. Kovács for useful discussions and a careful reading of the manuscript, and I. Varga for useful discussions. MG was partially supported by the NKFIH grant KKP-126769. This work was also supported by the NKFIH excellence Grant TKP2021-NKTA-64.